Journalist Report Feb 19th

Crew 205 Crew Journalist Report 19-FEB-2019 Sol 9

Author: Maria Grulich

Title: “What Goes Up!”

Mars awaited our rising out of our crew quarters with beautiful sunshine through our windows. The smell of coffee was around the table while we prepared our breakfast: mainly oats and cereals.

The meal plan was already made by Hannah, our Chef de Cuisine: chocolate lava cake for tomorrow and salmon soup for lunch. The cake for today will be baked by Maria, and it is a German apple pie.

The crew’s commander is Ghanim, and as every day, he led us as we planned what we want to do. Therefore, people spread to do their tasks. Nathan and Dave’s second home became the science dome as they analyzed the newly found shell fossils and separated the rock samples to decide which ones are worth to bring back to Earth.

Natalia is working on guidelines for future crews. Crew 205 would like to give some help for future emerging space leaders.

Gus, our mascot, is still very active helping explain information about Mars and following crew members around. Maria made some progress with video editing and hopes to send out first videos by tomorrow. Veronica opened the first Mars gym in the downstairs area of the Hab so that crew members could work off the hushpuppies and Hannah’s amazing cooking in the last days.

Everyone is catching up on surveys.

Before the comms window Mars surprised us with one of its brightest Moons since we got here, it is glorious. Tomorrow will be another day on Mars and we can’t wait to discover more amazing DVD and CDs of the almost endless collection here in the Hab, we leave you with this title: “What goes up” Space, Trance, Transfusion.

Tomorrow a geographically inspired lava chocolate cake will await us! Ooh la la!

Ad Astra!

Operations Report Feb 19th

Crew 205 Operations Report 19 – FEB – 2019

SOL: 9

Name of person filing report: Verónica Triviño

Non-nominal systems: Generator

Notes on non-nominal systems: Radiator hoses and alternator need replacement on generator. A contractor has been arranged to make these repairs within a few days.

Generator (hours run): Turned on at 1530 (19th Feb), turned off at 0730 (19th Feb)

Solar— SOC % – Turned off Gen (19th Feb) 95%, Turned on Gen (19th Feb) 57%

Diesel Reading – 75%

Propane Reading – ~ 36%

Ethanol Free Gasoline – 10 gallons

Water (Auxillary tank) – Not in use gallons

Water (Static tank) – around 46%, ~ 251 gallons

Auxillary to Static tank transfer – No

Gallons transferred: Not Applicable

Water in GreenHab – ~ 198 gallons

Water (loft) – 12 marked level

Static to Loft Pump used – yes

Water Meter: 01413601 units

Toilet tank emptied: No

Deimos rover used: No, still not functional

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not Applicable

Beginning charge: Not Applicable

Ending charge: Not Applicable

Currently charging: Not Applicable

Opportunity rover used: Not Used

Hours: 51.0 hours

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Curiosity rover used: Not Used

Hours: 75.0 hours

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Spirit rover used: Not Used

Hours: 75.7

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Notes on rovers: N/A

ATV’s Used: N/A

Reason for use: Not used

Oil Added? N/A

ATV Fuel Used: N/A

# Hours the ATVs were Used today: None

Notes on ATVs: None

HabCar used and why, where? N/A

Crew Car used and why, where? N/A

General notes and comments: Nothing to report

Summary of internet: Nothing to report

Summary of suits and radios: Nothing to report.

Summary of Hab operations: Nothing to report

Summary of Green Hab operations: Nothing to report

Summary of Science Dome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: We still need a 75W standard light bulb.

GreenHab Report – Feb 19th

Crew 205 Green Hab Report – 19-FEB-2019

GreenHab Officer: Nathan Hadland

Environmental control: Heating and Cooling w/ ambient air (1 hrs)

Shade Cloth (40%)

Average temperatures:
Low: 18.4°C

High: 31.7°C

Hours of supplemental light: 5 Hours

Daily water usage for crops: 7.5 gallons

Water in Blue Tank: ~ 198 gallons

Time(s) of watering for crops:

TIME NOTED TEMPERATURE(° C ) / HUMIDITY(%)
[AT TIME OF WATERING]

1) 09.30 AM 20.6° C / 27 %
2) 12.30 PM 31.4° C / 21 %
3) 15:30 PM 23.9° C / 22 %
4) 18:00 PM 23.5° C / 32 %

Changes to crops: More “Mystery Greens” began to sprout and sprouts were emerging around the planter.

Narrative: It is becoming clear that not all regolith simulants are created equal. Several Arabidopsis thaliana experiments are performing very well and others continue to dry out, likely due to the alkaline nature and wettability properties. In addition, the paper towel germinations have finished and we may transfer them to pots for future crews to maintain.

We are decreasing watering frequency due to the constant saturation of the soil and to allow the substrate to aerate. The door remained closed for most of the day due to the low ambient temperatures. However, when the environmental data was checked at noon, the temperature had jumped to 31 degrees C due to the intense sunlight. However, after approximately one hour of ambient airflow, the temperature had dropped to 19 degrees so the door was closed at 1330.

Harvest: None.

Support/supplies needed: None.

Sol Summary Report – Feb 19th

CREW 205 Sol Summary

Sol: 9

Sol Summary Title: A Martian Day in the Hab

Author: Ghanim Alotaibi

Mission status: Nominal

Sol activity summary: The bad weather last night affected today’s activities. EVA plans were cancelled last night, and astronomical observations submitted a few days ago needed to wait for an extra night. This morning the crew woke up with an exceptional white view of the Martian landscape from the Hab window. While having breakfast, the white bright snow was melting down, and as we approached lunch time, the muddy Martian land started to finally appear.

Duties in the Hab, and its facilities started just after finishing breakfast as planned last night. Dave, Nathan and Veronica went to the Science Dome to clean, measure and characterize the samples. In fact, those samples made the whole crew so excited, and was the main topic of today and last night. Our XO today, Hannah, prepared lunch for the crew with the help of Dan. Maria, continued to work hard for our outreach movie, this time, with the help of Natalia. Natalia started preparing a guideline document for future crews and discussed with the commander of the day, Ghanim.

In addition to the above work, the crew continued the normal daily maintenance, cleaning of the Hab and writing reports as usual. The only new duty was to sweep out the snow accumulated in the tunnel to avoid muddy walk way.

In the night, Hannah cooked us a delicious fried rice with chicken, and the crew prepared to watch the Apollo 13 Film.

Anomalies: EVA cancelled

Weather: partly cloudy

Reports to be filed: Sol Summary, Operations Report, Greenhab Report, Journalist Report.

Support Requested: None